OpenStudy (annie96):

Help. Identify all of the unknowns in the problem. There are 9 more children than parents in a room. There are 25 people in the room in all. How many children are there in the room? (Points : 1) the number of children and the number of parents in the room the number of children in the room the number of parents in the room the number of persons in the room

OpenStudy (anonymous):

Use a system of equation. Use variables x for children and y for adults. x = y + 9 x + y = 25 Use the substitution method to solve, substitute y + 9 for x in the last problem. (y + 9) +y =25 2y + 9 =25 subtract 9 from both sides. 2y = 16 divide by 2 y = 8, there are 8 adults. there must be 17 kids. it works because 8 + 9 is 17.

OpenStudy (anonymous):

please give me a medal hope this helps.